OpenStudy (anonymous):

If E and F are events with P(EnF)=0.1, P(E|F)=0.25, and P(F|E)=0.125, find (a) P(E) (b) P(F)

OpenStudy (unklerhaukus):

\[P(E\cap F)=0.1=\frac1{10}\]\[P(E|F)=0.25=\frac14\]\[P(F|E=0.125=\frac 18\]

OpenStudy (unklerhaukus):

is this right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

does the vertical bar '|' mean 'or'

OpenStudy (unklerhaukus):

and the '\(\cap\)' mean intersection ?

OpenStudy (zarkon):

'|' means 'Given'

OpenStudy (anonymous):

P(E) = P(EnF)/P(F/E) P(F) = P(EnF)/P(E/F)